18 years ago father s 3 timez his son's age.now father s 2 timez his son's age .What s the present age of son ?

Answer by Cromlix(4381)

Hi there,
father s 2 times his son's age
F = 2T
18 years ago father s 3 times his son's age.
F - 18 = 3(T - 18)
Multiply out
F - 18 = 3T - 54
Replace F with 2T from top equation
2T - 18 = 3T - 54
Collect like terms
2T - 3T = -54 + 18
-T = -36
T = 36
Current ages:
Father = 72 yrs old
Timez = 36 yrs old
Hope this helps:-)